OpenStudy (anonymous):

Charles Law: A sample of helium occupies 1.00 L at a temperature of 250. K. What volume would it occupy at 1500. K?

OpenStudy (frostbite):

Hello @LaurenEBlaga and welcome to \( \LARGE \sf \bbox[#40B9E9]{\color{white}{Open}}\bbox[#A8CE91]{\color{white}{Study}}\) I don't recall Charles law, but all the information you need is in the ideal gas equation. Lets assume we work with an idea gas, we then write the ideal gas equation as: \[\Large pV=nRT \to \frac{ pV }{ T }=nR\] The amount of substance \(n\) remains constant in a condition change, so does the gas constant, meaning we can set up the equation \[\Large \frac{ p _{0}V _{0} }{ T _{0} }=\frac{ p _{1}V _{1} }{ T _{1} }\] Lets assume that the condition chance happens under constant pressure, that is \(p_{0}=p_{1}\). We can then cancel the pressure and we get \[\Large \frac{ V_{0} }{ T _{0} }=\frac{ V _{1} }{ T _{1} }\]

OpenStudy (frostbite):

Insert your values into the equation and evaluate the new volume.